A friend of mine is giving a dinner party. His current wine supply includes 8 bottles of zinfandel, 11 of merlot, and 10 of cabernet (he only drinks red wine), all from different wineries.

(a) If he wants to serve 3 bottles of zinfandel and serving order is important, how many ways are there to do this?
ways

(b) If 6 bottles of wine are to be randomly selected from the 29 for serving, how many ways are there to do this?
ways

(c) If 6 bottles are randomly selected, how many ways are there to obtain two bottles of each variety?
ways

(d) If 6 bottles are randomly selected, what is the probability that this results in two bottles of each variety being chosen? (Round your answer to three decimal places.)


(e) If 6 bottles are randomly selected, what is the probability that all of them are the same variety? (Round your answer to three decimal places.)
Answer #1

Solution:-
a) n = 8 , k = 3

=> 8*7*6 = 336

b) Combination = 29c6 = 475020

permutation : 29*28*27*26*25*24 = 342014400

c) use the product for k tuples,her k = 3 is the number of types of wine : (8C2)(11C2)(10C2) = 28*55*45 = 69300

d) => 69300/475020 = 0.146

e) => (28+462+210)/475020 = 0.001